# Parallelizing the Approximate Minimum Degree Ordering Algorithm: Strategies and Evaluation

### v1

@@
Abstract.
The approximate minimum degree algorithm is widely used before numerical factorization to reduce fill-in for sparse matrices. While considerable attention has been given to the numerical factorization process, less focus has been placed on parallelizing the approximate minimum degree algorithm itself. In this paper, we explore different parallelization strategies, and introduce a novel parallel framework that leverages multiple elimination on distance-2 independent sets. Our evaluation shows that parallelism within individual elimination steps is limited due to low computational workload and significant memory contention. In contrast, our proposed framework overcomes these challenges by parallelizing the work across elimination steps. To the best of our knowledge, our implementation is the first scalable shared memory implementation of the approximate minimum degree algorithm. Experimental results show that we achieve up to an 8.30x speedup using 64 threads over the state-of-the-art sequential implementation in SuiteSparse.
1. Introduction
Efficiently solving sparse linear systems is a cornerstone of many scientific and engineering applications. Cholesky factorization, a widely used direct method for solving symmetric positive definite systems, often encounters challenges when the given matrix $A$ is sparse. In such cases, zero entries in $A$ can transform into nonzeros, known as fill-ins, in the Cholesky factor. These fill-ins can significantly increase both memory usage and computational cost, undermining the efficiency of the factorization. To address this, fill-reducing orderings seek to find a permutation matrix $P$ such that the permuted matrix $PAP^{T}=LL^{T}$ reduces the number of fill-ins in the Cholesky factor $L$. However, minimizing the number of fill-ins is an NP-complete problem (Yannakakis, 1981), and as a result, various heuristics have been developed, such as nested dissection (George, 1973; Gilbert and Tarjan, 1986) and approximate minimum degree (AMD) (Amestoy et al., 1996). This paper solely focuses on parallelizing the AMD algorithm in shared memory since to the best of our knowledge, no parallel implementation currently exists.
The need for a scalable shared memory parallel AMD implementation is motivated by literature and practical use cases. Trotter et al. (Trotter et al., 2023) recently conducted an extensive study on 490 matrices and found that while nested dissection and the AMD algorithm achieved similar levels of fill reduction, the ordering time for nested dissection was up to an order of magnitude longer than that of AMD. Furthermore, while significant work has been done to parallelize nested dissection on both shared and distributed memory systems (Chevalier and Pellegrini, 2008; Karypis and Kumar, 1998; LaSalle and Karypis, 2015), no similar efforts have been made for the AMD algorithm, highlighting a key gap in the literature. Meanwhile, GPU Cholesky solvers have steadily improved over time, as shown in Table 1, driven by advances in software, and are expected to continue accelerating with future GPU generations. In contrast, the standalone sequential AMD implementation on CPU has emerged as a critical bottleneck. Our goal in this paper is therefore to design a scalable shared memory parallel AMD implementation, closing the gap in literature and in practice.
Table 1. Time (in seconds) for the sequential AMD implementation from SuiteSparse (v7.10.1) (Davis, 2019) on an AMD EPYC 7763 CPU, and for Cholesky solver cuSolverSp (Corporation, 2025b) (v11.6.1) and its successor cuDSS (Corporation, 2025a) (v0.5.0) to solve the reordered symmetric positive definite system on an A100 GPU. The matrices are randomly permuted symmetrically beforehand to decouple tie-breaking issues in the AMD algorithm.
Matrix Name SuiteSparse AMD cuSolverSp cuDSS
nd24k 0.83 117.17 4.11
ldoor 1.28 11.01 3.55
Flan_1565 4.18 out of memory 16.85
Parallelizing the AMD algorithm presents two main challenges. First, each elimination step involves selecting a pivot vertex with the minimum approximate degree, eliminating it, and then updating its neighborhood and the approximate degrees of the neighbors. These steps are inherently sequential, as a new pivot cannot be selected until all updates from the previous pivot are complete. While neighborhood and approximate degree updates could, in theory, be parallelized, the opportunities for efficient parallelism are limited due to little amount of work and high memory contention. Second, although the multiple elimination strategy from the sequential multiple minimum degree algorithm (Liu, 1985) can help reduce dependencies by selecting a maximal independent set of pivots, this approach introduces significant overlap in these pivots’ neighborhoods. While such overlap is beneficial in the sequential setting—reducing the cost of degree updates—it introduces contention on the underlying graph structure and complicates the maintenance of approximate degrees, ultimately hindering scalability when attempting to parallelize the AMD algorithm using multiple elimination.
In this paper, we address these challenges by selecting distance-2 independent sets of pivots, enabling parallel elimination steps while keeping approximate degree updates simple and efficient. We further exploit key properties of the AMD algorithm to design specialized concurrent data structures—including a concurrent graph representation and concurrent approximate degree lists—that eliminate the need for dynamic memory allocation and significantly reduce memory contention. Together, these optimizations lead to the first scalable shared memory implementation of the AMD algorithm, outperforming the long-standing sequential version in SuiteSparse (Davis, 2019). Our contributions include:
  • We present the first scalable shared memory implementation of the AMD algorithm. Our approach leverages multiple elimination on distance-2 independent sets to extract parallelism across elimination steps, and introduces specialized concurrent data structures tailored to the unique properties of the AMD algorithm.
  • Our implementation achieves ordering quality comparable to the state-of-the-art, with tunable control via relaxation and limitation factors to balance fill-in and parallelism.
  • We demonstrate up to an 8.30×8.30$\times$8.30 × speedup using 64 threads on a diverse set of real-world sparse matrices compared to the leading sequential implementation in SuiteSparse. Our in-depth performance analysis highlights current bottlenecks and points to promising directions for future research.
2. Preliminaries
Before introducing the approximate minimum degree (AMD) algorithm (Amestoy et al., 1996), we begin with its predecessors, the minimum degree algorithm (Rose, 1972) and the multiple minimum degree algorithm (Liu, 1985), to establish the foundational concepts. From this point onward, we disregard the positive definiteness of the matrices, as it does not impact the ordering algorithms discussed in this paper.
Figure 1. An example illustrating how elimination graphs work. For demonstration purposes, we eliminate vertices 5, 2, and 9 in order, rather than following the minimum degree criterion. When a vertex is eliminated, its neighbors form a clique.
@@ 2. Preliminaries
An example to demonstrate how elimination graphs work.
Figure 2. An example illustrating how quotient graphs work. For demonstration purposes, we eliminate variables 5, 2, and 9 in order, rather than following the minimum degree criterion. Circles represent variables and squares represent elements.
@@
An example to demonstrate how quotient graphs work.
2.1. Minimum Degree Algorithm
The minimum degree algorithm is most naturally described using elimination graphs (Rose, 1972). Let $A$ be an $n\times n$ symmetric sparse matrix. The sparsity pattern of $A$ can be represented by a graph $G^{0}=(V^{0},E^{0})$, where $V^{0}=\{1,2,\cdots,n\}$, and each vertex $i$ corresponds to row/column $i$ of $A$. An edge $(i,j)$ is in $E^{0}$ if and only if $A_{i,j}\neq 0$. The elimination graph $G^{k}=(V^{k},E^{k})$ represents the sparsity pattern of the remaining submatrix after the first $k$ pivots have been selected and eliminated during Cholesky factorization. Let $N^{k}_{v}$ denote the neighborhood of vertex $v$ in the elimination graph $G^{k}$. The elimination process begins with $G^{0}$. To obtain $G^{k}$ from $G^{k-1}$, the $k$th pivot $p$ is selected from $V^{k-1}$, removed from the vertex set, and its neighbors are connected to form a clique—reflecting the fill-ins introduced by an outer product update in Cholesky factorization. An example is presented in Figure 1. Formally, the update can be expressed as:
@@ 2.1. Minimum Degree Algorithm
The minimum degree algorithm selects, at each step, the vertex with the smallest degree in $G^{k-1}$ as the $k$th pivot. Since multiple vertices can have the same degree, a tie-breaking strategy is required. Ties are typically resolved arbitrarily or based on the initial ordering; however, it has been shown that the choice of tie-breaking strategy can significantly affect the number of fill-ins produced (George and Liu, 1989).
2.2. Quotient Graphs
A key challenge with elimination graphs is that the storage needed for $G^{k}$ is not known in advance due to the formation of cliques during elimination, which necessitates dynamic memory management. The quotient graph (George and Liu, 1980), also known as the generalized element model or superelement (Eisenstat et al., 1976; Speelpenning, 1978; Duff and Reid, 1983), addresses this issue by representing cliques more compactly—storing only the set of vertices in each clique. George and Liu (George and Liu, 1981) showed that the quotient graph takes no more storage than the original graph $G^{0}$.
Following historical terminology, we refer to eliminated vertices from the original graph as elements, and those yet to be eliminated as variables. To stay consistent with the literature, we adopt the notations used in (Amestoy et al., 1996). The quotient graph $\mathcal{G}^{k}=\left(V^{k},\overline{V}^{k},\mathcal{A}^{k},\mathcal{E}^{k},\mathcal{L}^{k}\right)$ provides an implicit representation of the elimination graph $G^{k}$. Here, $V^{k}$ denotes the set of variables, $\overline{V}^{k}$ the set of elements, $\mathcal{A}^{k}$ the set of variable-to-variable connections, $\mathcal{E}^{k}$ the connections from variables to elements, and $\mathcal{L}^{k}$ the connections from elements to variables. Notably, there is no connection between elements, and while $\mathcal{E}^{k}$ and $\mathcal{L}^{k}$ are essentially transposes of each other, they are often maintained separately in practical implementations for efficiency. In essence, the variables $V^{k}$ correspond to the uneliminated vertices, with $\mathcal{A}^{k}$ being the remaining edges from the original graph; the elements $\overline{V}^{k}$ represent cliques formed during elimination, with their connections compactly encoded in $\mathcal{E}^{k}$ and $\mathcal{L}^{k}$.
@@ 2.2. Quotient Graphs
$\mathcal{A}_{v}$ =$=\{v^{\prime}\mid(v,v^{\prime})\in\mathcal{A}\}\subseteq V,$
$\mathcal{E}_{v}$ =$=\{e^{\prime}\mid(v,e^{\prime})\in\mathcal{E}\}\subseteq\overline{V},$
$\mathcal{L}_{e}$ =$=\{v^{\prime}\mid(e,v^{\prime})\in\mathcal{L}\}\subseteq V.$
The quotient graph and the elimination graph are closely related through the following invariant: for any vertex $v$ in $G$, which also appears as a variable in $\mathcal{G}$, its neighborhood in the elimination graph can be recovered by
(1) $N_{v}=\left(\mathcal{A}_{v}\cup\bigcup_{e\in\mathcal{E}_{v}}\mathcal{L}_{e}\right)\setminus\{v\}.$
In other words, the neighborhood of $v$ in the elimination graph is reconstructed from the quotient graph by interpreting the remaining original edges and the clique information encoded in the adjacent elements.
When a variable $p$ is selected as the $k$th pivot from $\mathcal{G}^{k-1}$, it becomes an element in $\mathcal{G}^{k}$. Its neighborhood $N_{p}$ is computed using Equation (1). To preserve the invariant for the remaining variables, the clique induced by $p$, namely $\{p\}\times N_{p}$, is added to $\mathcal{L}$, and similarly $N_{p}\times\{p\}$ is added to $\mathcal{E}$. Only the set of variables forming the clique is stored, instead of their pairwise connections. Additionally, connections now covered by this newly formed clique are removed. Formally, the connection updates are defined as:
@@
Figure 2 illustrates how quotient graphs work, where circles represent variables and squares represent elements. Taking $\mathcal{G}^{1}$ as an example, we determine $N_{2}=\{1,3,4,6,8\}$ based on Equation (1) using the information encoded in element 5 and the original edges. This result is consistent with $G^{1}$ shown in Figure 1. Once variable 2 is eliminated, a new adjacency set $\mathcal{L}_{2}=N_{2}$ is formed. Redundant variable-variable connections, such as the connection between 1 and 4, are removed in the new quotient graph $\mathcal{G}^{2}$, as they can be implied from $\mathcal{L}_{2}$. Additionally, element 5 is removed since the clique it encoded is now covered by the new clique represented by element 2 using $\mathcal{L}_{2}$.
@@
To build the minimum degree algorithm using quotient graphs, it is necessary to update the degrees of all variables $v\in N_{p}$ after eliminating the pivot $p$. However, this step is computationally expensive because reconstructing the neighborhood of each $v$ requires performing set union operations, as described in Equation (1). The cost of these operations is linear in the total size of the sets involved. These degree updates have been identified as the primary bottleneck of the minimum degree algorithm in the literature (Liu, 1985; George and Liu, 1989).
2.3. Multiple Minimum Degree Algorithm
@@
The multiple minimum degree algorithm (Liu, 1985) addresses the bottleneck in degree updates by selecting a maximal independent set of pivots, denoted by $M$, within an additive relaxation of the minimum degree criterion in the elimination graph. All pivots in $M$ are eliminated sequentially with their associated connection updates applied before any degree updates are performed on affected variables. The independence of these pivots ensures that pivots do not interfere with one another, while the maximality of the set increases the likelihood of overlapping pivots’ neighborhoods. This overlap is key to reducing the cost of degree updates: whereas the standard minimum degree algorithm requires a separate degree update for each adjacent pivot in $M$, the multiple minimum degree algorithm defers degree updates until all pivots in $M$ are eliminated, allowing each affected variable to perform only a single, consolidated degree update.
@@
For example, consider selecting a maximal independent set $M=\{1,3,7,9\}$ from $\mathcal{G}^{0}$ in Figure 2. If the variables in $M$ are eliminated one by one using the minimum degree algorithm, then each of the variables 2, 4, 6, and 8 will incur two degree updates. In contrast, with multiple elimination, where all variables in $M$ are eliminated in a single step, only one degree update is required for each of the variables 2, 4, 6, and 8 after all connection updates are applied.
Figure 3. An example illustrating how to compute approximate degrees after eliminating variable 7. To determine the exact degree of variable 8 in $\mathcal{G}^{4}$, which is 4, one would need to compute $|\mathcal{L}_{2}\cup\mathcal{L}_{7}\cup\mathcal{L}_{9}\setminus\{8\}|$. A naïve estimate using a union bound results in double counting, as variables 4 and 6 appear twice, yielding an estimate of 6. The approximate degree, however, mitigates such double counting by leveraging information already captured in $\mathcal{L}_{7}$ (from the pivot 7). Specifically, the estimate is computed as $|\mathcal{L}_{7}\setminus\{8\}|+|\mathcal{L}_{2}\setminus\mathcal{L}_{7}|+|\mathcal{L}_{9}\setminus\mathcal{L}_{7}|=5$, with only variable 6 being double counted.
[…]
An example to demonstrate how approximate minimum degrees work.
2.4. Approximate Minimum Degree Algorithm
@@
The approximate minimum degree (AMD) algorithm (Amestoy et al., 1996) adopts an alternative strategy to the multiple minimum degree algorithm for reducing the cost of degree updates. Instead of computing exact degrees for all variables $v$ in $\mathcal{L}_{p}=N^{k-1}_{p}$ after the connection updates, the AMD algorithm estimates an upper bound, $d^{k}_{v}$, for the exact degree:
$d^{k}_{v}=\min\left\{\begin{array}[]{l}n-k-1,\\ d^{k-1}_{v}+|\mathcal{L}_{p}\setminus\{v\}|-1,\\ |\mathcal{A}_{v}\setminus\{v\}|+|\mathcal{L}_{p}\setminus\{v\}|+\sum_{e\in\mathcal{E}_{v}}|\mathcal{L}_{e}\setminus\mathcal{L}_{p}|\end{array}\right\}$
@@
with $d^{0}_{v}$ initialized as the exact degree of $v$ in the original graph. Note that $d_{v}$ remains unchanged if $v\notin\mathcal{L}_{p}$. This estimate is known as the approximate degree. The first term reflects the maximum possible degree given the size of the remaining submatrix, the second accounts for the worst case fill-in introduced by $p$, and the third refines the estimate by avoiding double counting using local information from the neighborhood of $p$. An example is shown in Figure 3.
@@
To efficiently compute $|\mathcal{L}_{e}\setminus\mathcal{L}_{p}|$, Algorithm 1 uses the identity $|\mathcal{L}_{e}\setminus\mathcal{L}_{p}|=|\mathcal{L}_{e}|-|\mathcal{L}_{e}\cap\mathcal{L}_{p}|$. During the scan, if an element $e$ is encountered for the first time, $w(e)$ is initialized to $|\mathcal{L}_{e}|$, and then decremented once for every variable $v\in\mathcal{L}_{e}\cap\mathcal{L}_{p}$. Otherwise, $w(e)$ remains negative. Combining both cases, the final value of $|\mathcal{L}_{e}\setminus\mathcal{L}_{p}|$ for each $e\in\overline{V}$ is given by (Amestoy et al., 1996):
@@
Algorithm 1 Computation of $|\mathcal{L}_{e}\setminus\mathcal{L}_{p}|$ for all $e\in\overline{V}$ (Amestoy et al., 1996)
[…]
2:$w(e)\leftarrow|\mathcal{L}_{e}|$EndIf
[…]
@@
By maintaining approximate degrees, the AMD algorithm selects, at each step, the variable with the smallest approximate degree in $\mathcal{G}^{k-1}$ as the $k$th pivot. This approach significantly reduces the cost of degree updates, as it avoids explicitly accessing the adjacency sets $\mathcal{L}_{e}$, which are required for exact degree computation in Equation (1) for the minimum degree algorithm. For a detailed complexity analysis of both the AMD and the minimum degree algorithms, we refer the reader to (Heggernes et al., 2002).
@@
The AMD algorithm also incorporates several additional techniques, including mass elimination (George and McIntyre, 1978), element absorption (Duff and Reid, 1983), indistinguishable nodes (George and Liu, 1981), and external degrees (Liu, 1985). These techniques are included in our parallel implementation of the AMD algorithm, though we omit detailed discussion here, as they are orthogonal to parallelization and may be explored in the cited references for interested readers.
2.5. Related Work
2.5.1. Minimum Degree and its Variants
@@
The minimum degree heuristic, a symmetric analog to Markowitz’s algorithm (Markowitz, 1957), was first proposed by Tinney and Walker (Tinney and Walker, 1967). A graph-theoretical interpretation was later developed by Rose (Rose, 1972), who coined the term minimum degree algorithm, as it selected the pivot with the smallest degree at each elimination step. Over the years, numerous enhancements have been proposed and have become standard components in modern implementations. These include mass elimination (George and McIntyre, 1978), indistinguishable nodes (George and Liu, 1981), quotient graphs (George and Liu, 1980) (also known as generalized element models or superelements (Eisenstat et al., 1976; Speelpenning, 1978; Duff and Reid, 1983)), incomplete degree updates (Liu, 1985; Eisenstat et al., 1982), element absorption (Duff and Reid, 1983), multiple elimination (Liu, 1985), and external degrees (Liu, 1985). These techniques and their impacts were comprehensively summarized by George and Liu (George and Liu, 1989).
2.5.2. Approximate Minimum Degree and its Variants

### v2

@@
Abstract
@@
1 Introduction.
Efficiently solving sparse linear systems is a cornerstone of many scientific and engineering applications. Cholesky factorization, a widely used direct method for solving symmetric positive definite systems, often encounters challenges when the given matrix $A$ is sparse. In such cases, zeros in $A$ can transform into nonzeros, known as fill-ins, in the Cholesky factor. These fill-ins can significantly increase both memory usage and computational cost, undermining the efficiency of the factorization. To address this, fill-reducing orderings seek to find a permutation matrix $P$ such that the permuted matrix $PAP^{T}=LL^{T}$ reduces the number of fill-ins in the Cholesky factor $L$. However, minimizing the number of fill-ins is an NP-complete problem [yannakakis1981computing], and as a result, various heuristics have been developed, such as nested dissection [george1973nested, gilbert1986analysis] and approximate minimum degree (AMD) [amestoy1996approximate]. This paper solely focuses on parallelizing the AMD algorithm in shared memory since to the best of our knowledge, no parallel implementation currently exists.
The need for a scalable shared memory parallel AMD implementation is motivated by literature and practical use cases. Trotter et al. [trotter2023bringing] recently conducted an extensive study on 490 matrices and found that while nested dissection holds some advantage over AMD in reducing fill-ins, the ordering time for nested dissection is up to an order of magnitude longer than that of AMD. Furthermore, while significant work has been done to parallelize nested dissection [chevalier2008pt, karypis1998parallel, lasalle2015efficient], no similar efforts have been made for the AMD algorithm, highlighting a key gap in the literature. Meanwhile, as shown in Table 1.1, GPU Cholesky solvers have steadily improved over time. In contrast, the standalone sequential AMD implementation on CPU is poised to become a bottleneck as other components of the solver continue to improve. To make the AMD algorithm more future-proof, our goal is therefore to design a scalable shared memory AMD implementation, closing the gap in theory and in practice.
Table 1.1: Time (in seconds) for the sequential AMD from SuiteSparse (v7.12.1) [suitesparsegithub] on an AMD EPYC 7763 CPU, and for Cholesky solver cuSolverSp [cusolver] (v11.6.1) and its successor cuDSS [cudss] (v0.7.1) to solve the reordered system on an A100 GPU in double precision.
Matrix AMD cuSolverSp cuDSS
nd24k 0.82 117.17 1.97
ldoor 1.42 11.01 3.03
Flan_1565 4.85 out of memory 18.92
Cube5317k 13.94 out of memory 43.90
Parallelizing the AMD algorithm presents two main challenges. First, each elimination step involves selecting a pivot vertex with minimum approximate degree, eliminating it, and then updating its neighborhood and the approximate degrees of the neighbors. These steps are inherently sequential, as a new pivot cannot be selected until all updates from the previous pivot are complete. While neighborhood and approximate degree updates could, in theory, be parallelized, the opportunities for efficient parallelism are limited due to little amount of work and high memory contention. Second, although the multiple elimination strategy from the sequential multiple minimum degree algorithm [liu1985modification] can help reduce dependencies by selecting a maximal independent set of pivots, this approach introduces significant overlap in these pivots’ neighborhoods. While such overlap is beneficial in the sequential setting, it introduces contention on the underlying graph structure and complicates the maintenance of approximate degrees, ultimately hindering scalability when attempting to parallelize the AMD algorithm using multiple elimination.
@@
2 Preliminaries.
Figure 2.1: An example illustrating how elimination graphs work. For demonstration purposes, we eliminate vertices 5, 2, and 9 in order, rather than following the minimum degree criterion. When a vertex is eliminated, its neighbors form a clique.
Figure 2.2: An example illustrating how quotient graphs work. For demonstration purposes, we eliminate variables 5, 2, and 9 in order, rather than following the minimum degree criterion. Circles represent variables and squares represent elements. When variable 2 is eliminated in $\mathcal{G}^{1}$, its neighborhood in $G^{1}$ is recovered using the clique encoded on element 5 and the remaining original neighbors, forming a new clique on variables 1, 3, 4, 6, and 8.
Before reviewing the approximate minimum degree (AMD) algorithm [amestoy1996approximate], we begin with its predecessors, the minimum degree algorithm [rose1972graph] and the multiple minimum degree algorithm [liu1985modification]. From this point onward, we disregard the positive definiteness of the matrices, as it does not impact the ordering algorithms discussed in this paper.
Table 2.1: Summary of notation used.
Symbol Description
AA An $n\times n$ sparse symmetric matrix
$p$ The selected pivot
$G^{k}$ The $k$th elimination graph
$\mathcal{G}^{k}$ The $k$th quotient graph
$V^{k}$ The set of vertices in $G^{k}$, equivalent to
the set of variables in $\mathcal{G}^{k}$
$E^{k}$ The set of edges in $G^{k}$
$\overline{V}^{k}$ The set of elements in $\mathcal{G}^{k}$
$N^{k}_{v}$ The neighborhood of vertex $v$ in $G^{k}$
$\mathcal{A}^{k}$ The connections among $V^{k}$ in $\mathcal{G}^{k}$
$\mathcal{E}^{k}$ The connections from $V^{k}$ to $\overline{V}^{k}$ in $\mathcal{G}^{k}$
$\mathcal{L}^{k}$ The connections from $\overline{V}^{k}$ to $V^{k}$ in $\mathcal{G}^{k}$
$\mathcal{A}^{k}_{v}$ The variables adjacent to variable $v$ in $\mathcal{G}^{k}$
$\mathcal{E}^{k}_{v}$ The elements adjacent to variable $v$ in $\mathcal{G}^{k}$
$\mathcal{L}^{k}_{e}$ The variables adjacent to element $e$ in $\mathcal{G}^{k}$
2.1 Minimum Degree Algorithm.
The minimum degree algorithm is most naturally described using elimination graphs [rose1972graph]. Let $A$ be an $n\times n$ symmetric sparse matrix. The sparsity pattern of $A$ can be represented by a graph $G^{0}=(V^{0},E^{0})$, where $V^{0}=\{1,2,\cdots,n\}$, and each vertex $i$ corresponds to row/column $i$ of $A$. An edge $(i,j)$ is in $E^{0}$ if and only if $A_{i,j}\neq 0$. The elimination graph $G^{k}=(V^{k},E^{k})$ represents the sparsity pattern of the remaining submatrix after the first $k$ pivots have been selected and eliminated during Cholesky factorization. Let $N^{k}_{v}$ denote the neighborhood of vertex $v$ in the elimination graph $G^{k}$. The elimination process begins with $G^{0}$. To obtain $G^{k}$ from $G^{k-1}$, the $k$th pivot $p$ is selected from $V^{k-1}$, removed from the vertex set, and its neighbors are connected to form a clique—reflecting the fill-ins introduced by an outer product update in Cholesky factorization. An example is presented in Figure 2.1. Formally, the update can be expressed as:
@@
The minimum degree algorithm selects, at each step, the vertex with the smallest degree in $G^{k-1}$ as the $k$th pivot. Since multiple vertices can have the same degree, a tie-breaking strategy is required. Ties are typically resolved arbitrarily or based on the initial ordering; however, it has been shown that the choice of tie-breaking strategy can significantly affect the number of fill-ins produced [george1989evolution].
2.2 Quotient Graphs.
A key challenge with elimination graphs is that the storage needed for $G^{k}$ is not known in advance due to the formation of cliques during elimination, which necessitates dynamic memory management. The quotient graph [george1980fast], also known as the generalized element model or superelement [eisenstat1976applications, speelpenning1978generalized, duff1983multifrontal], addresses this issue by representing cliques more compactly—storing only the set of vertices in each clique. George and Liu [george1981computer] showed that the quotient graph takes no more storage than the original graph.
Following historical terminology, we refer to eliminated vertices from the original graph as elements, and those yet to be eliminated as variables. To stay consistent with the literature, we adopt the notations used in [amestoy1996approximate]. The quotient graph $\mathcal{G}^{k}=\left(V^{k},\overline{V}^{k},\mathcal{A}^{k},\mathcal{E}^{k},\mathcal{L}^{k}\right)$ provides an implicit representation of the elimination graph $G^{k}$. Here, $V^{k}$ denotes the set of variables, $\overline{V}^{k}$ the set of elements, $\mathcal{A}^{k}$ the set of variable-to-variable connections, $\mathcal{E}^{k}$ the connections from variables to elements, and $\mathcal{L}^{k}$ the connections from elements to variables. Notably, there is no connection between elements, and while $\mathcal{E}^{k}$ and $\mathcal{L}^{k}$ are essentially transposes of each other, they are maintained separately in practical implementations for efficiency. In essence, the variables $V^{k}$ correspond to the uneliminated vertices, with $\mathcal{A}^{k}$ being the remaining edges from the original graph; the elements $\overline{V}^{k}$ represent cliques formed during elimination, with their connections compactly encoded in $\mathcal{E}^{k}$ and $\mathcal{L}^{k}$.
@@
$\mathcal{A}_{v}$ $=\{v^{\prime}\mid(v,v^{\prime})\in\mathcal{A}\}\subseteq V,$
$\mathcal{E}_{v}$ $=\{e^{\prime}\mid(v,e^{\prime})\in\mathcal{E}\}\subseteq\overline{V},$
$\mathcal{L}_{e}$ $=\{v^{\prime}\mid(e,v^{\prime})\in\mathcal{L}\}\subseteq V.$
The quotient graph and the elimination graph are closely related through the following invariant: for any vertex $v$ in $G$, which also appears as a variable in $\mathcal{G}$, its neighborhood is:
(2.1) $N_{v}=\left(\mathcal{A}_{v}\cup\bigcup_{e\in\mathcal{E}_{v}}\mathcal{L}_{e}\right)\setminus\{v\}.$
In other words, the neighborhood of $v$ in the elimination graph can be reconstructed from the quotient graph by interpreting the remaining original neighbors and the clique information encoded in the adjacent elements.
When a variable $p$ is selected as the $k$th pivot from $\mathcal{G}^{k-1}$, it becomes an element in $\mathcal{G}^{k}$. Its neighborhood $N_{p}$ is computed using Equation (2.1). To preserve the invariant for the remaining variables, the new clique induced by $p$ is added to $\mathcal{L}$ encoded as $\{p\}\times N_{p}$, and similarly $N_{p}\times\{p\}$ is added to $\mathcal{E}$. Additionally, connections now covered by this newly formed clique are removed. An example is presented in Figure 2.2. Formally, the connection updates are defined as:
@@
To build the minimum degree algorithm using quotient graphs, it is necessary to update the degrees of all variables $v\in N_{p}$ after eliminating the pivot $p$. However, this step is computationally expensive because reconstructing the neighborhood of each $v$ requires performing set union operations, as described in Equation (2.1). These degree updates have been identified as the primary bottleneck of the minimum degree algorithm in the literature [liu1985modification, george1989evolution].
2.3 Multiple Minimum Degree Algorithm.
@@
The multiple minimum degree algorithm [liu1985modification] addresses the bottleneck in degree updates by selecting a maximal independent set of pivots, denoted by $M$, within an additive relaxation of the minimum degree criterion. All pivots in $M$ are eliminated sequentially with their associated connection updates applied before any degree updates are performed on affected variables. The independence of these pivots ensures that pivots do not interfere with one another, while the maximality of the set increases the likelihood of overlapping pivots’ neighborhoods. This overlap is key to reducing the cost of degree updates: whereas the standard minimum degree algorithm requires a separate degree update for each adjacent pivot in $M$, the multiple minimum degree algorithm defers degree updates until all pivots in $M$ are eliminated, allowing each affected variable to perform only a single, consolidated degree update.
@@
For example, consider selecting a maximal independent set $M=\{1,3,7,9\}$ from $\mathcal{G}^{0}$ in Figure 2.2. If the variables in $M$ are eliminated one by one using the minimum degree algorithm, then each of the variables 2, 4, 6, and 8 will incur two degree updates. In contrast, with multiple elimination, where all variables in $M$ are eliminated in a single step, only one degree update is required for each of the variables 2, 4, 6, and 8 after all connection updates are applied.
Figure 2.3: An example illustrating how to compute approximate degrees after eliminating variable 7. To get the approximate degree of variable 8 in $\mathcal{G}^{4}$, we compute $|\mathcal{L}_{7}\setminus\{8\}|+|\mathcal{L}_{2}\setminus\mathcal{L}_{7}|+|\mathcal{L}_{9}\setminus\mathcal{L}_{7}|=5$.
2.4 Approximate Minimum Degree Algorithm.
@@
The approximate minimum degree (AMD) algorithm [amestoy1996approximate] adopts an alternative strategy to reduce the cost of degree updates. Instead of computing exact degrees, the AMD algorithm estimates an upper bound, $d^{k}_{v}$, for the exact degree of a vertex $v$ in $G^{k}$:
$d^{k}_{v}=\min\left\{\hskip-3.00003pt\begin{array}[]{l}n-k-1,\\ d^{k-1}_{v}+|\mathcal{L}_{p}\setminus\{v\}|-1,\\ |\mathcal{A}_{v}\setminus\{v\}|+|\mathcal{L}_{p}\setminus\{v\}|+\sum_{e\in\mathcal{E}_{v}}|\mathcal{L}_{e}\setminus\mathcal{L}_{p}|\end{array}\hskip-3.00003pt\right\}$
@@
with $d^{0}_{v}$ initialized as the exact degree of $v$ in $G^{0}$. Note that $d_{v}$ remains unchanged if $v\notin\mathcal{L}_{p}$. This estimate is known as the approximate degree. The first term reflects the maximum possible degree given the size of the remaining submatrix, the second accounts for the worst case fill-in introduced by $p$, and the third refines the estimate using union bound and avoids double counting using local information from the neighborhood of $p$. An example is shown in Figure 2.3.
@@
To efficiently compute $|\mathcal{L}_{e}\setminus\mathcal{L}_{p}|$, Algorithm 1 uses the identity $|\mathcal{L}_{e}\setminus\mathcal{L}_{p}|=|\mathcal{L}_{e}|-|\mathcal{L}_{e}\cap\mathcal{L}_{p}|$. During the scan, if an element $e$ is encountered for the first time, $w(e)$ is set to $|\mathcal{L}_{e}|$, and then decremented once for every variable $v\in\mathcal{L}_{e}\cap\mathcal{L}_{p}$. Otherwise, $w(e)$ remains negative. The negativity assumption on $w(e)$ required by the algorithm is ensured using a timestamp method, as detailed in [amestoy1996approximate]. Combining both cases, the final value of $|\mathcal{L}_{e}\setminus\mathcal{L}_{p}|$ for each $e\in\overline{V}$ is given by [amestoy1996approximate]:
@@
Algorithm 1 Computation of $|\mathcal{L}_{e}\setminus\mathcal{L}_{p}|$ [amestoy1996approximate]
1: Assume $w(e)<0$ for all $e\in\overline{V}$.
2: for each variable $v\in\mathcal{L}_{p}$ do
3:   for each element $e\in\mathcal{E}_{v}$ do
4:    if $w(e)<0$ then
5:      $w(e)\leftarrow|\mathcal{L}_{e}|$
6:    $w(e)\leftarrow w(e)-1$
@@
The AMD algorithm therefore selects, at each step, the variable with the smallest approximate degree in $\mathcal{G}^{k-1}$ as the $k$th pivot. For a detailed complexity analysis of both the AMD and the minimum degree algorithms, we refer the readers to [heggernes2002computational].
@@
The AMD algorithm also incorporates several additional techniques, including mass elimination [george1978application], element absorption [duff1983multifrontal], indistinguishable nodes [george1981computer], and external degrees [liu1985modification]. These techniques are included in our parallel implementation of the AMD algorithm, though we omit detailed discussion here, as they are orthogonal to parallelization and may be explored in the cited references for interested readers.
2.5 Related Work.
2.5.1 Minimum Degree and its Variants.
@@
The minimum degree heuristic, a symmetric analog to Markowitz’s algorithm [markowitz1957elimination], was first proposed by Tinney and Walker [tinney1967direct]. A graph-theoretical interpretation was later developed by Rose [rose1972graph], who coined the term minimum degree algorithm. Over the years, numerous enhancements have been proposed and have become standard components in modern implementations. These include mass elimination [george1978application], indistinguishable nodes [george1981computer], quotient graphs [george1980fast] (also known as generalized element models or superelements [eisenstat1976applications, speelpenning1978generalized, duff1983multifrontal]), incomplete degree updates [liu1985modification, eisenstat1982yale], element absorption [duff1983multifrontal], multiple elimination [liu1985modification], and external degrees [liu1985modification]. These techniques and their impacts were comprehensively summarized by George and Liu [george1989evolution].
2.5.2 Approximate Minimum Degree and its Variants.